OpenStudy (leozap1):

Fan & Medal, multiple choice. Javier owes $4,289 on a credit card with a 17.5% interest rate. What is the monthly payment he should make to pay off this debt in 12 months, assuming he does not charge any more purchases with the card? A. $71.43 B. $357.42 C. $392.20 D. $877.24

OpenStudy (leozap1):

I know it's not A or B. I can eliminate them because A is too low of a payment, and B would be the payment without interest.

OpenStudy (leozap1):

I'm thinking the answer is C.

OpenStudy (greatlife44):

Well, i'm assuming that that the interest rate is monthly? So i think you would need a way to represent the total amount of money made in 12 months 17.5% so 4289 is the base fee, and each month you've got interest factored in. \[P(x) = 4289+1.175x \] this will give you the total number of money he has to pay in 12 months adjusted for interest.

OpenStudy (greatlife44):

\[P(x) = 4289+1.175(12) = (total-payment)\] \[\frac{ total~payment }{ 12 } = monthly~payment\]

OpenStudy (leozap1):

Okay, so P(x)=4303.1?

OpenStudy (leozap1):

Divided by 12 that gives me 358.5917 (rounded)

OpenStudy (leozap1):

It doesn't say if it's monthly or anything so I'm not sure if it's monthly.

OpenStudy (greatlife44):

P(12) = 4303.1/12 = 358 remember it's in dollars :)

OpenStudy (leozap1):

So it would be $358.59? What do I do from here though, as it's not one of my answer choices lol.

OpenStudy (greatlife44):

Pretty sure it's going to be 357, I may have not rounded properly along somewhere.